OpenStudy (anonymous):

Bacteria in a culture are growing exponentially with time, as shown in the table below. Day Bacteria 0 100 1 200 2 400 Which of the following equations expresses the number of bacteria, y, present at any time, t? the answers will be posted below

OpenStudy (anonymous):

I think you might want to ask this in Biology.... I'll take a look anyways.

OpenStudy (anonymous):

y=100 + 2^t \[y=(100)\times(2)^t\] y=2^t \[y=(200) \times (2)^t\]

OpenStudy (anonymous):

these are the choices

OpenStudy (anonymous):

Option B without a doubt.

OpenStudy (anonymous):

so how do we solve this? or at least figure out the answer?

OpenStudy (anonymous):

What I did was that I took the basic equation to be A(x^m) From that one skeletal equation I just took the values.

OpenStudy (anonymous):

wait so what do you mean by plugging in the values?

OpenStudy (anonymous):

Yes. It's pretty much similar to the question you previously asked (first order reaction)

OpenStudy (anonymous):

umm?

OpenStudy (anonymous):

is there like an equation that you can show me how to set up?

OpenStudy (anonymous):

whats the formula again?